OpenStudy (anonymous):

Guys I give medals The side of a square is measured to be 10 ft with a possible error of ±0.1 ft. Use differentials to estimate the error in the calculated area

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

A = s^2

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@satellite73 @satellite73 @satellite73 @satellite73 @satellite73

OpenStudy (anonymous):

@poo

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

@satellite73 @satellite73

OpenStudy (anonymous):

@Gokuporter

jimthompson5910 (jim_thompson5910):

A = s^2 dA/ds = 2s dA = 2s*ds plug in the given side length (s = 10) and the error in the side length measurement (ds = 0.1) to find the error in the area measurement (dA)